OpenStudy (study_buddy99):

Alg. 2?! evaluate: 81 3/4

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle (81)^{3/4} }\) ?

OpenStudy (study_buddy99):

there are no parenthesis around it

OpenStudy (solomonzelman):

No parenthesis around 81? That is all the same.

OpenStudy (solomonzelman):

Another rule you need to know: \(\color{#000000}{ \displaystyle \large \color{blue}{a}^{\color{green}{b}/\color{red}{c}}=\sqrt[\Large \color{red}{c}]{\color{blue}{a}^\color{green}{b}} }\)

OpenStudy (study_buddy99):

I'm still slightly confused so is the answer: